OpenStudy (anonymous):

what is 21y times y?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

21y*y=21(y*y)=21y^2

OpenStudy (anonymous):

21y^2

OpenStudy (anonymous):

21y^2

OpenStudy (anonymous):

Thanks :))